• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name folts

    Log entry time 16:31:36 on April 08, 2007

    Entry number 198236

    keyword=spectrometer move

    Moved left spectrometer to ~140 degrees. Turned off power to bogies.

    Will re-enter Hall tomorrow to move back to forward position.

    NOTE: This is a high rad experiment and we are on the edge of needing a RWP to enter the target area. There are high rad areas in the Hall!